Circular No. 8466

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 B1 (CHRISTENSEN) E. J. Christensen reports his discovery of a comet (discovery observa-

tion below; 10′′ tail in p.a. 250^{o}) on images taken with the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey. Following posting on the ‘NEO Confirmation Page’, J. Young reported that his CCD images taken through moderate cirrus clouds with the Table Mountain 0.6-m reflector on Jan. 16.5 UT show a 5′′ coma and a broad tail 25′′ long spanning p.a. 240^{o}–260^{o}. M. Tichý found the object to be diffuse with a tail in p.a. 255^{o} on images taken by himself and J. Tichá with the Kleť 1.06-m telescope.

2005 UT  α_{2000}  δ_{2000}  Mag.
Jan. 16.46322  14^{h}28^{m}27.^{s}36  +19^{o}00′31′′.6  17.4

The available astrometry, the following very uncertain parabolic orbital elements, and an ephemeris appear on MPEC 2005-B13.

T = 2005 Feb. 14.399 TT  ω = 34.^{o}624
Ω = 192.641 }2000.0
q = 5.28914 AU  i = 86.060

COMETS C/2004 W7–W11, X4–X9 (SOHO) Further to IAUC 8458, following are the initial observations of addi-

tional comets found on SOHO website images (XG = X. Gao; RM = R. Matson; SF = S. Farmer, Jr.); all are Kreutz sungrazers except for C/2004 W10 (which belongs to the Marsden group) and C/2004 X7 (which does not belong to a known group).

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 W7 Nov. 24.096 15^{h}51.^{m}3 −24^{o}34′ C3/2 TH 2005-A12
C/2004 W8  24.129 15 59.8 −22 18 C2  XL 2005-A12
C/2004 W9  28.435 16 18.7 −23 07 C2  XG 2005-A12
C/2004 W10  29.102 16 18.4 −21 20 C2  RK 2005-A13
C/2004 W11  30.618 16 28.8 −23 32 C2  RM 2005-A13
C/2004 X4  Dec. 2.971 16 35.9 −25 46 C3/2 SF 2005-A13
C/2004 X5  4.976 16 50.6 −24 19 C2  TH 2005-A13
C/2004 X6  6.643 16 58.2 −24 30 C2  TH 2005-A40
C/2004 X7  7.185 17 01.4 −21 35 C2  RK 2005-A40
C/2004 X8  8.811 17 08.2 −24 47 C2  RM 2005-A40
C/2004 X9  13.242 17 27.7 −25 04 C2  KB 2005-A40
2005 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green